OpenStudy (anonymous):

Please correct me if I go wrong at any point. Let A be a matrix. A is symmetric if A = transpose(A) A is orthogonal if transpose(A) = inverse(A) A is orthogonal AND symmetric implies A = transpose(A) = inverse(A) Correct?

OpenStudy (anonymous):

Then A is normal if \[AA ^{T} = A ^{T}A\] But how can it be implied that A is normal if A is either symmetric or orthogonal? (Looking for a short proof I feel like I just need a small leap in my logic to wrap my head around this)

OpenStudy (anonymous):

hmm i think i prooved the orthogonal case: considering a seperate property that is true for all invertible matrices \[AA ^{-1} = A ^{-1}A = I\] so if \[A ^{T} = A ^{-1}\] (definition of orthogonal matrix A) this implies \[AA ^{T} = A ^{T}A = I\] proves that an orthogonal matrix is also normal Still trying to work out why a symmetric matrix is also normal

OpenStudy (anonymous):

If A is symmetric, then we have:\[AA^T=(A)A^T=(A^T)A^T=A^T(A^T)=A^T(A)=A^TA\]I know it looks weird, i added in more parenthesis to emphasis where I was using the fact that A^T = A since A is symmetric.